OpenStudy (anonymous):

What is the solution to the system of equations shown below? 2x-y+3z=8 x-6y-z=0 -6x+3y-9z=24

OpenStudy (anonymous):

Do you want the solution or the method on how to do it?

OpenStudy (anonymous):

Let me just explain it real fast to you. Since these equations are made of the same variables you can add them. So when you subtract 2x(eq2) from eq(1) you get: 11y +5z= 8. Continuing this will get you the answers you need.

OpenStudy (anonymous):

sorry I still need help

OpenStudy (anonymous):

have you studied matrices and/or determinants?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

ok. multiply the first eq. with "3" and add it to the third eq. what do you see?

OpenStudy (anonymous):

wait why are we multiplying a 3?

OpenStudy (anonymous):

you will realize a hard fact about this system :)

OpenStudy (anonymous):

ok. so when multiplying a 3 to the first eq. I get 6x-3y+9z=24 -6x+3y-9z=24 wait so they cancel out? where do we go from there?

OpenStudy (anonymous):

exactly. this is what is called as an "under-determined" system. there are infinite solutions possible.

OpenStudy (anonymous):

you have "3" unknown variables and essentially, only "2" unique relations.

OpenStudy (anonymous):

where does the second equation come into play?

OpenStudy (anonymous):

for a system of "n" variables, you will need ATLEAST "n" unique equations.

OpenStudy (anonymous):

?

OpenStudy (anonymous):

well, we'd have used it if there were "3" unique equations like before (the Kirchhoff's laws problem)

OpenStudy (anonymous):

but since we have established that there is no "unique" solution, we cannot solve it.

OpenStudy (anonymous):

why wouldn't it be no solution?

OpenStudy (anonymous):

but you can go about this way... \[\text{let}\qquad z=a\\ 2x-y=8-3a\qquad\text{and}\qquad x-6y=a\\ \;\\ -12x+6y=-48+18a\\ \oplus\\ x-6y=a\\ ---------------\\ -11x=-48+9a\implies\boxed{x={48\over11}-{18\over11}a} \]

OpenStudy (anonymous):

you can try the rest of your life, if you've got that time, to try to come up with a unique solution but you wont . so, lets get our basics set up before we loose our minds :)

OpenStudy (anonymous):

so its infinitely many, opposed to no solution because? sorry

OpenStudy (anonymous):

because it is an "under-determined" system. (just out of curiosity, double check the equations that you wrote here with the problem, are the exactly same?)

OpenStudy (anonymous):

yes. They are the same

OpenStudy (anonymous):

ok. then there you go... like I explained, there are "3" unknowns and only two unique equations. so, the system is under-determined and no unique solution exists.